# Gravitational Bound State Perturbations Inside Black Holes and Isospectrality

### v1

@@ Abstract
We study the bound state solutions for the polar perturbations in the interior of the Schwarzschild black hole. It is shown that for a given value of the spherical harmonic index $\ell$, there are a total of $\ell-1$ bound states for polar perturbations. We show both analytically and numerically that the spectrum of $\ell-2$ of these perturbations agrees identically with the spectrum of axial perturbations. Correspondingly, the isospectrality between the bound states of the axial and polar perturbations in the interior of the black hole does hold. This is significant since the effective potential is singular at the center of the black hole, and the existence of the isospectrality is far from obvious. Furthermore, we demonstrate that the additional mode found in the spectrum of polar perturbations is the algebraically special mode, which is the ground state of polar perturbations as well.
@@
The quasi-normal modes (QNMs) perturbations in the exterior of BHs are vastly studied, for some reviews and an incomplete list of theoretical studies see Nollert (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Berti et al. (2025); Iyer and Will (1987); Motl and Neitzke (2003); Andersson and Howls (2004); Cardoso et al. (2003); Medved et al. (2004); Berti et al. (2006); Maggiore (2008); Konoplya et al. (2019); Hod (2025). The scalar, vector, and the axial tensor perturbations are governed by the Regge-Wheeler (RW) equation, while the corresponding equation for the polar tensor perturbations is the Zerilli equation Regge and Wheeler (1957); Zerilli (1970); Chandrasekhar (1985). As the Zerilli potential is more complicated than the RW potential, the bound state perturbations of polar perturbations were not studied in Steinhauer et al. (2025) and Firouzjahi (2025). In the exterior region of BHs, there is a one-to-one transformation which links the polar perturbations to the axial perturbations Chandrasekhar (1985); Chandrasekhar and Detweiler (1975); Anderson and Price (1991). Since this transformation is regular in the exterior region, an immediate conclusion is that the axial and polar perturbations have the same spectra of QNMs perturbations.
In this work, we extend upon the analysis of Steinhauer et al. (2025); Firouzjahi (2025) and study the bound state perturbations for the polar perturbations in the interior of the Schwarzschild BH. It is important to note that the bound state solutions considered here and in Firouzjahi (2018); Steinhauer et al. (2025); Firouzjahi (2025) are different from the bound state solutions of the inverted potential proposed by Mashhoon and collaborators Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996). More specifically, the bound states in Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996) are obtained from the complex transformation of the RW equation to bring the potential in the form of Poschl-Teller or the Eckart potentials which can be solved analytically. This method was further employed recently in Völkel (2025) (see also Hatsuda (2020)) in which the inverted RW potential is considered directly and a numerical method is used to determine the eigenvalues of energies. In contrast, the bound states in our study are interpreted as the perturbations in the interior of the BH with imaginary frequencies such that their profiles are regular and confined to the interior of the BH which are inaccessible to an exterior observer. With these regularity boundary conditions, the profiles of the bound states are square integrable in the interior of the BH.
@@
The specific forms of $V^{\pm}$ are presented in Appendix A while their plots are shown in Fig. 1. It is seen that $V^{-}$ develops a negative global minimum before diverging like $V^{-}\sim 3/r^{4}$ near the center. On the other hand, $V^{+}$ is always negative and diverges like $V^{+}\sim-1/r^{4}$ at $r=0$. The fact that $V^{-}$ develops a negative global minimum indicates that the bound states with $\omega=i\,\omega_{I}$ can exist for axial perturbations Steinhauer et al. (2025); Firouzjahi (2025) . However, the existence of bound states for $V^{+}$ may look non-trivial. Indeed, we confirm that the bound states for polar perturbations exist as well.
@@
The ASM is the solution in which the left and right hand sides of Eq. (4) vanish simultaneously. With $\omega_{I}>0$, the left hand side of Eq. (4) allows only the lower sign associated to the polar perturbations with the spectrum Chandrasekhar (1985, 1984); Couch and Newman (1973); Wald (1973),
$\hskip 28.45274pt\omega_{\mathrm{sp}}=\frac{i}{6}(\ell-1)\ell(\ell+1)(\ell+2)=\frac{2i}{3}\lambda(1+\lambda)\,.$ (6)
@@
in which $N_{+}(\lambda)$ is a normalization constant. One can easily check that $Z^{+}_{\mathrm{sp}}(r)$ satisfies the Zerilli equation (2). Of course, the second order differential equation (2) has another solution for the polar perturbations. But, constructing the second solution via the Wronskian method, we have checked specifically that this second solution is singular at the horizon $r=1$ and is not allowed as a consistent bound state solution. On the other hand, since $Z^{+}_{\mathrm{sp}}(r)$ satisfies Eq. (4) identically at $\omega=\omega_{\mathrm{sp}}$, there is no equation left for $Z^{-}_{\mathrm{sp}}(r)=0$, meaning that the ASM is not shared by the axial perturbations. This is consistent with the results of Firouzjahi (2025) in which no bound state solutions for axial perturbations at the frequency $\omega=\omega_{\mathrm{sp}}$ was obtained.
In the next section we confirm that for all other bound states (i.e. $\omega\neq\omega_{\mathrm{sp}}$) the transformations (3) and (4) are regular. An immediate conclusion is that the isospectrality between the axial and polar bound state perturbations does hold. We confirm this interesting result by direct numerical analysis as well.
@@
As mentioned before, the equation $\hat{\mathcal{D}}^{\pm}_{\ell}\,\phi^{\pm}_{\ell,0}=0$ specifies the ground states of these potentials with $\mathcal{E}^{\pm}_{\ell,0}=0$. Correspondingly, the ground states associated to each Hamiltonian are obtained to be,
$\phi^{+}_{\ell,0}(r)\sim\dfrac{1}{\phi^{-}_{\ell,0}(r)}\sim\dfrac{r\,(r-1)^{-i\omega_{\mathrm{sp}}}}{(3+2\lambda r)}\,e^{-i\omega_{\mathrm{sp}}\,r}\,.$ (13)
The ground state $\phi^{+}_{\ell,0}(r)$ is regular and normalizable in the domain of the superpotential while the state $\phi^{-}_{\ell,0}$, is singular at $r=0,1$ and non-normalizable; this confirm that the SUSY is unbroken.
As $U^{\pm}_{\ell}$ are partner potentials and SUSY is not broken, the isospectrality between their spectra does hold Gangopadhyaya et al. (2017). However, the ground state of $U^{+}_{\ell}$ can not be obtained from any of the eigenfunction of $U^{-}_{\ell}$ by operation of $\hat{\mathcal{D}}_{\ell}^{-}$, and it does not have any correspondence in the spectrum of $U^{-}_{\ell}$.
@@
As mentioned, the Schrodinger-like equation (12) is equivalent to the original equation (2). Consequently, the eigenfunctions of polar perturbations are given by $Z^{+}_{\ell,n}=\phi^{+}_{\ell,n}$. More specifically, the ground state is given by $Z^{+}_{\ell,n=0}(r)=\phi^{+}_{\ell,0}(r)$ with $\phi^{+}_{\ell,0}(r)$ as given in Eq. (13) with the spectrum $\omega^{+}_{\ell,n=0}=\omega_{\mathrm{sp}}$. Furthermore, the eigenfunctions of axial perturbations are given by $Z^{-}_{\ell,n}=\phi^{-}_{\ell,n}$. Finally, from Eqs. (14) and (15) and the fact that there are only $\ell-2$ axial bound states, we obtain,
@@
Numerical Results: Our numerical results also confirm the isospectrality and the existence of the ASMs. We present the detail numerical analysis in the Appendix A. Below is the summary of our main numerical results.
(a): For a given value of $\ell$, there are total $\ell-1$ bound state solutions for the polar perturbations. This should be compared with the case of axial perturbations, which admit $\ell-2$ solutions Firouzjahi (2025). (b): All $\ell-2$ spectra of the axial perturbations are identically shared with the polar perturbations. In this view, the isospectrality between the polar and axial perturbations does hold. We have confirmed this conclusion to 40 digits in our numerical analysis. (c): The additional mode in polar perturbations is indeed the ASM with the spectrum given in Eq. (6). This is the ground state solution for the polar perturbations in which, for a given value of $\ell$, $\omega^{2}$ has the least value (i.e., the most negative value). In particular, for $\ell=2$, $\omega_{\mathrm{sp}}$ is the only allowed spectrum. We have confirmed that the spectrum of ASM agrees with Eq. (6) to 40 digits. This is significant, noting that there are controversies on the exact numerical values of ASMs for QNMs Berti et al. (2009); Maassen van den Brink (2000); Leaver (1985); Nollert (1993); Liu and Mashhoon (1996). (d): The level number of the excited states with $n=1,2,$... is the same as the number of the nodes for the profile of $Z(r)$. For example, the first excited state has one node, the second excited state has two nodes, and so on. On the other hand, the most excited state with $n=\ell-2$ has $\ell-2$ nodes. (e): For large $\ell$ and for low-lying states (say $n=1$), the spectrum satisfies an analytic WKB formula as given in Firouzjahi (2025) in which $\omega_{I}\propto(\ell+\frac{1}{2})^{4}$. (f): For $\ell\gg 1$, the most excited state, which for a given value of $\ell$ has the lowest value of $\omega_{I}$, satisfies $\omega_{I}=1+c\,\ell^{-1}+{\cal O}(\ell^{-2})$ with $c\simeq 1$. An important conclusion from the above results is that independent of $\ell$ and $n$, all modes satisfy the lower bound $\omega_{I}>1$. Note that the conclusions (d), (e) and (f) listed above are identical to axial perturbations obtained in Firouzjahi (2025). Furthermore, conclusions (b) and (c) are confirmed analytically from our SUSY analysis as reviewed in previous section.
@@
In Table 1, we have presented $\omega_{I}$ for some values of $\ell$. We have calculated the spectrum in 40 digits but in order to save space, here we only present the spectrum to 10 digits. One can check that, except for $\omega_{\mathrm{sp}}$ associated to the ground state $n=0$, all the remaining $\omega_{I}$ are the same as for axial perturbations presented in Firouzjahi (2025).
@@
Finally, it is helpful to examine the similarities and differences of our bound states with QNMs in the exterior region. In both cases, the isospectrality is valid and the polar perturbations have the additional ASM excitations. However, a key difference is that for a given value of $\ell$ the number of bound states in the interior of the BH is finite but the number of QNMs is infinite. This is perhaps due to differences between the asymptotic shapes of the corresponding effective potentials in the interior and exterior regions. More specifically, $V^{-}_{\mathrm{eff}}$ can be approximated by Coulomb potential in far regions Liu and Mashhoon (1996). However, near the horizon in the interior $(r_{*}\rightarrow-\infty)$ it falls off exponentially like $e^{r_{*}}$ Firouzjahi (2018) (see Appendix A). Nonetheless, the question of why the number of bound states in the interior of BH is finite (for a given value of $\ell$) is important which deserves careful investigations. Another difference is that ASM in our analysis is the ground state ($n=0$), while in the spectrum of QNMs the ASM is far from the ground state. For example, for $\ell=2,3$, ASM corresponds to $n=9$ and $n=41$ in QNMs spectrum respectively Leaver (1985).
@@
Conclusions: In this work, we have studied the bound states of the polar perturbations in the interior of the Schwarzschild BH and examined the isospectrality. The bound states are solutions with imaginary spectrum which are regular at the center of BH while their profiles fall off exponentially near the event horizon.
@@
We have shown numerically that for a given value of $\ell$, there are $\ell-1$ bound states of polar perturbations in which the spectra of $\ell-2$ of these bound states are identical (within 40 digits in our numerical analysis) with those of axial perturbations. In this sense, the isospectrality between these two perturbations does hold. However, there is an additional mode in the spectrum of polar perturbations which turns out to be the ASM with the spectrum Eq. (6) and the mode function given in Eq. (7). The isospectrality between the polar and axial perturbations is a direct consequence of the map that links these two perturbations as we studied via SUSY method. The fact that the potentials are singular at $r=0$ does not destroy this correspondence. Unlike the QNMs spectrum, the number of bound states (for a given value of $\ell)$ is finite and the ASM is the ground state. As in Firouzjahi (2025), the bound states satisfy the lower bound $\omega>1$ which is saturated only asymptotically for $\ell\rightarrow\infty$. In Tables 1 we have presented the numerical results for $\omega_{I}$ for some values of $\ell$ while the profiles of the mode functions are presented in Figs. 2 and 3. From Fig. 3 we observe that the number of nodes is equal to the level number $n$.

### v2

@@ Abstract
We study the bound state solutions for the polar perturbations in the interior of the Schwarzschild black hole. It is shown that for a given value of the spherical harmonic index $\ell$, there are a total of $\ell-1$ bound states for polar perturbations. We show both analytically and numerically that the spectrum of $\ell-2$ of these perturbations coincides exactly with the spectrum of axial perturbations. Consequently, the isospectrality between the bound states of axial and polar perturbations in the interior of the black hole is preserved. Furthermore, the additional mode found in the spectrum of polar perturbations is the algebraically special mode, which also furnishes the ground state of polar perturbations. It is shown that the spectrum of the highly excited states is equally spaced, which, in the semi-classical approximation, yields the black hole area quantization $\Delta A=16\pi l_{\mathrm{Pl}}^{2}$.
@@
The quasi-normal modes (QNMs) perturbations in the exterior of BHs are vastly studied, for some reviews and an incomplete list of theoretical studies see Nollert (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Berti et al. (2025); Iyer and Will (1987); Motl and Neitzke (2003); Andersson and Howls (2004); Cardoso et al. (2003); Medved et al. (2004); Berti et al. (2006); Konoplya et al. (2019); Hod (2025). The scalar, vector, and the axial tensor perturbations are governed by the Regge-Wheeler (RW) equation, while the corresponding equation for the polar tensor perturbations is the Zerilli equation Regge and Wheeler (1957); Zerilli (1970); Chandrasekhar (1985). The RW and Zerilli potentials have different algebraic forms, but there is a one-to-one transformation which links the polar perturbations to the axial perturbations Chandrasekhar (1985); Chandrasekhar and Detweiler (1975); Anderson and Price (1991). Since this transformation is regular in the exterior region, an immediate conclusion is that the axial and polar perturbations have the same spectra of QNMs perturbations.
In this work, we extend upon the analysis of Steinhauer et al. (2025); Firouzjahi (2025) and study the bound state perturbations for the polar perturbations in the interior of the Schwarzschild BH. It is important to note that the bound state solutions considered here and in Firouzjahi (2018); Steinhauer et al. (2025); Firouzjahi (2025) are different from the bound state solutions of the inverted potential proposed by Mashhoon and collaborators Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996). More specifically, the bound states in Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996) are obtained from the complex transformation of the RW equation to bring the potential in the form of Poschl-Teller or the Eckart potentials which can be solved analytically. This method was further employed recently in Völkel (2025) (see also Hatsuda (2020)) in which the inverted RW potential is considered directly and a numerical method is used to determine the eigenvalues of energies.
@@
The specific forms of $V^{\pm}$ are presented in Appendix A while their plots are shown in Fig. 1. It is seen that $V^{-}$ develops a negative global minimum before diverging like $V^{-}\sim 3/r^{4}$ near the center. On the other hand, $V^{+}$ is always negative and diverges like $V^{+}\sim-1/r^{4}$ at $r=0$.
@@
The ASM is the solution in which the left and right hand sides of Eq. (4) vanish simultaneously. With $\omega_{I}>0$, the left hand side of Eq. (4) allows only the lower sign corresponding to the polar perturbations with the spectrum Chandrasekhar (1985, 1984); Couch and Newman (1973); Wald (1973),
$\hskip 14.22636pt\omega=\omega_{\mathrm{sp}}\equiv\frac{i}{6}(\ell-1)\ell(\ell+1)(\ell+2)=\frac{2i}{3}\lambda(1+\lambda)\,.$ (6)
@@
in which $N_{+}(\lambda)$ is a normalization constant. One can easily check that $Z^{+}_{\mathrm{sp}}(r)$ satisfies the Zerilli equation (2). On the other hand, since $Z^{+}_{\mathrm{sp}}(r)$ satisfies Eq. (4) identically at $\omega=\omega_{\mathrm{sp}}$, there is no equation left for $Z^{-}_{\mathrm{sp}}(r)=0$, meaning that the ASM is not shared by the axial perturbations. This is consistent with the results of Firouzjahi (2025) in which no bound state solutions for axial perturbations at the frequency $\omega=\omega_{\mathrm{sp}}$ was obtained.
@@
As mentioned, the Schrodinger-like equation (12) is equivalent to the original equation (2). Consequently, the eigenfunctions of polar perturbations are given by $Z^{+}_{\ell,n}=\phi^{+}_{\ell,n}$. In particular, the ground state is given by $Z^{+}_{\ell,n=0}(r)=\phi^{+}_{\ell,0}(r)$, with $\phi^{+}_{\ell,0}(r)$ as given in Eq. (7) and with the spectrum $\omega^{+}_{\ell,n=0}=\omega_{\mathrm{sp}}$. Furthermore, the eigenfunctions of axial perturbations are given by $Z^{-}_{\ell,n}=\phi^{-}_{\ell,n}$. Finally, from Eqs. (13) and (14) and the fact that there are only $\ell-2$ axial bound states, we obtain,
@@
Numerical Results and Implications: Our numerical results also confirm the isospectrality and the existence of the ASMs. We present the detail numerical analysis in the Appendix A. Below is the summary of our main numerical results and their physical implications.
(a): For a given value of $\ell$, there are total $\ell-1$ bound state solutions for the polar perturbations. This should be compared with the case of axial perturbations, which admit $\ell-2$ solutions Firouzjahi (2025). (b): All $\ell-2$ spectra of the axial perturbations are identically shared with the polar perturbations. In this view, the isospectrality between the polar and axial perturbations does hold. We have confirmed this conclusion to 40 digits in our numerical analysis. (c): The additional mode in polar perturbations is indeed the ASM with the spectrum given in Eq. (6). This is the ground state solution for the polar perturbations in which, for a given value of $\ell$, $\omega^{2}$ has the least value (i.e., the most negative value). In particular, for $\ell=2$, $\omega_{\mathrm{sp}}$ is the only allowed spectrum. We have confirmed that the spectrum of ASM agrees with Eq. (6) to 40 digits. This is significant, noting that there are controversies on the exact numerical values of ASMs for QNMs Berti et al. (2009); Maassen van den Brink (2000); Leaver (1985); Nollert (1993); Liu and Mashhoon (1996). (d): The level number of the excited states with $n=1,2,...,\ell-2$ is the same as the number of the nodes in the profile of $Z(r)$. For example, the first excited state has one node, the second excited state has two nodes, and so on. On the other hand, the most excited state with $n=\ell-2$ has $\ell-2$ nodes. (e): For large $\ell$ and for low-lying states (say $n=1$), the spectrum satisfies an analytic WKB formula as given in Firouzjahi (2025) in which $\omega_{I}\propto(\ell+\frac{1}{2})^{4}$. (f): For $\ell\rightarrow\infty$, the spectrum of most excited states with $n=\ell-2,\ell-3,$..., are equally spaced with $\omega_{I}=(\ell-n)-1$. An important conclusion from the above results is that independent of $\ell$ and $n$, all modes satisfy the lower bound $\omega_{I}>1$ with the most excited state having the minimum value of $\omega_{I}$ Firouzjahi (2025). Note that the conclusions (d), (e) and (f) listed above are identical to axial perturbations obtained in Firouzjahi (2025). Furthermore, conclusions (b) and (c) are confirmed analytically from our SUSY analysis as reviewed in previous section.
@@
In Table 1, we have presented $\omega_{I}$ for some values of $\ell$. We have calculated the spectrum in 40 digits but in order to save space, here we only present the spectrum to 10 digits. One can check that, except for $\omega_{\mathrm{sp}}$ associated to the ground state $n=0$, all the remaining $\omega_{I}$ are the same as in axial perturbations presented in Firouzjahi (2025).
@@
As mentioned in point (f) listed above, the spectrum of perturbations for the most excited states $n=\ell-2,\ell-3,$... have the important universal property that $\omega_{I}$ are equally spaced for $\ell\rightarrow\infty$. This was shown for $s=0,1,2$ in Firouzjahi (2025), and following the same analysis, one can check that this is true for $s=1/2,3/2$ as well. This has important implications for the understanding of the BH area quantization as follows.
Following the logic of Hod (1998); Maggiore (2008), one can use Bohr’s correspondence principle for the transition between two states of $n,n^{\prime}\gg 1$ to obtain a semi-classical expression for Bekenstein’s BH area quantization Bekenstein (1974). For this prescription to be applicable, the spectrum should depend only on the BH intrinsic properties and not on the properties of the fields under consideration. Indeed, the bound states in the interior of BHs satisfy this universal property in which, independent of $s$, $(2GM)\,\omega_{I}(n)\rightarrow(\ell-n)-1$ for $n\rightarrow\infty$ (here we have restored the scaling $2GM$). Now consider the transition from the most excited state $n=\ell-2$ to the second most excited state $n^{\prime}=\ell-3$. This yields $(2GM)\Delta\omega=1$, inducing a change in the mass of BH, $\Delta M=\hbar\Delta\omega=\hbar(2GM)^{-1}$. On the other hand, from the BH area formula $A=16\pi G^{2}M^{2}$ we have $\Delta A=32\pi G^{2}M\Delta M$. Using our predictions for $\Delta M$, this yields the BH area quantization,
$\hskip 14.22636pt\Delta A=32\pi G^{2}M\Delta M=16\pi l_{\mathrm{Pl}}^{2}\,,\quad\quad(l_{\mathrm{Pl}}^{2}\equiv\hbar G)\,.$ (16)
In comparison, Maggiore Maggiore (2008) has used the transition between highly excited states of QNMs in the exterior region and obtained $\Delta A=8\pi l_{\mathrm{Pl}}^{2}$. Indeed, parameterizing the BH area quantization via $\Delta A=\alpha l_{\mathrm{Pl}}^{2}$, there are various estimations for the numerical factor $\alpha$ ranging from $4\ln 2$ to $32\pi$ Bekenstein (1974); Bekenstein and Mukhanov (1995); Mukhanov (1986); Hod (1998); Maggiore (2008); Coates et al. (2022) while our prediction is $\alpha=16\pi$.
@@
Finally, it is helpful to examine the similarities and differences of our bound states with QNMs in the exterior region. In both cases, the isospectrality is valid and the polar perturbations have the additional ASM excitations. However, a key difference is that for a given value of $\ell$ the number of bound states in the interior of the BH is finite but the number of QNMs is infinite. This is due to differences between the asymptotic shapes of the corresponding effective potentials in the interior and exterior regions. More specifically, $V^{-}_{\mathrm{eff}}$ can be approximated by the Coulomb potential in far regions Liu and Mashhoon (1996). However, near the horizon in the interior $(r_{*}\rightarrow-\infty)$ it falls off exponentially like $e^{r_{*}}$ Firouzjahi (2018) (see Appendix A). Another difference is that ASM in our analysis is the ground state ($n=0$), while in the spectrum of QNMs the ASM is far from the ground state. For example, for $\ell=2,3$, ASM corresponds to $n=9$ and $n=41$ in QNMs spectrum respectively Leaver (1985).
@@
Conclusions: In this work, we have studied the bound states of the polar perturbations in the interior of the Schwarzschild BH and examined various of its physical implications. The bound states are solutions with imaginary spectrum which are regular at $r=0$ while their profiles fall off exponentially near the event horizon.
@@
We have shown numerically that for a given value of $\ell$, there are $\ell-1$ bound states of polar perturbations in which the spectra of $\ell-2$ of these bound states are identical (within 40 digits in our numerical analysis) with those of axial perturbations. As a result, the isospectrality between these two perturbations does hold. However, there is an additional mode in the spectrum of polar perturbations which turns out to be the ASM with the spectrum Eq. (6) and the mode function Eq. (7). The isospectrality between the polar and axial perturbations is a direct consequence of the map that links these two perturbations as we studied via SUSY method. The fact that the potentials are singular at $r=0$ does not destroy this correspondence. Unlike the QNMs spectrum, the number of bound states is finite (for a given value of $\ell)$, with ASM being the ground state. In Tables 1 we have presented the numerical results for $\omega_{I}$ for some values of $\ell$ while the profiles of the mode functions are presented in Figs. 2 and 3. From Fig. 3 we observe that the number of nodes is equal to the level number $n$. As in Firouzjahi (2025), and independent of $s$, the bound states associated to the highly excited states are equally spaced. Employing this universal property, the BH area quantization in semi-classical approximation is calculated to be $\Delta A=16\pi l_{\mathrm{Pl}}^{2}$.